# Mask Image Watermarking

### v2

@@ 3.3 Inference
The inference process consists of three main stages: watermark embedding, localization, and extraction. Embedding is performed by the encoder. In MaskMark-D, the encoder embeds the watermark bits across the entire image. In MaskMark-ED, users provide both the watermark bits and a mask, allowing the encoder to embed the watermark into specific regions. Localization and extraction are performed by the decoder. As shown in Figure 2, unlike conventional global extraction methods, our approach first uses the decoder’s localization module to identify watermark-containing regions. The rest of the image is masked out by setting non-watermarked areas to zero, reducing noise during extraction. The watermark is then recovered from the retained regions. As discussed in Sec. 4.5, this strategy improves extraction robustness, especially when only small regions remain watermarked.
@@ 3.4 Usage Scenarios
This section outlines the usage scenarios of MaskMark-D and MaskMark-ED, which are tailored to different protection requirements. MaskMark-D is designed for full-image protection. It enables reliable watermark extraction even when parts of the image are tampered with, making it suitable for scenarios that demand global content integrity and tamper detection, such as copyright enforcement or image provenance verification. MaskMark-ED, on the other hand, focuses on protecting specific regions of interest, such as faces, logos, or sensitive content. It allows targeted verification or tracing if these regions are reused or misappropriated, without introducing watermarks across the entire image. This strategy not only supports privacy-aware or content-specific protection, but also tends to provide better robustness within the marked regions.
@@ 4.1 Implementation Details
For all experiments, we train MaskMark on 83k images from the MS-COCO 2014 training set [12] and the training details are provided in Appendix C.1.1.
@@
To ensure fair comparison, we fix the image resolution to $512\times 512$. For baseline methods that do not support this resolution, we follow the resolution scaling strategy from TrustMark [2] to interpolate the watermark strength (see Appendix C.2.1), which has been shown to preserve watermarking performance. To ensure comparable visual fidelity across variants, we set the JND scaling factor $\mu$ to 1.3 for MaskMark-D and 1.75 for MaskMark-ED. For robustness evaluation, we separately assess valuemetric and geometric distortions. For valuemetric robustness, we randomly sample from a set of ten common distortions, including JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. For geometric robustness, we randomly sample from three typical transformations: Rotation, Perspective, and Horizontal Flip. These distortions collectively cover the vast majority of real-world transformations that watermarked images are likely to encounter in practical scenarios. Detailed parameter settings for each distortion are provided in Appendix C.2.2. We also evaluate robustness under each specific distortion individually and the detailed results are presented in Appendix D.2.
@@ 4.2 Global and Local Watermarking Comparison
We compare MaskMark with seven recent open-source watermarking methods, including both global (e.g., StegaStamp [21], SepMark [23], TrustMark [2], EditGuard [28], Robust-Wide [7], VINE [13]) and local (WAM [19]) approaches. We use the clean (EditGuard-C) and degraded (EditGuard-D) variants of EditGuard, and the robust version (VINE-R) of VINE. For global watermarking, we evaluate on 1k images from the MS-COCO 2014 validation set using PSNR, SSIM, and Bit Accuracy, where PSNR and SSIM measure the visual quality of watermarked images, and Bit Accuracy evaluates watermark extraction performance. For local watermarking, we use all 41k validation images and evaluate Bit Accuracy under different watermarked area ratios. By default, MaskMark embeds 32 bits for fair comparison with WAM, though MaskMark supports flexible bit lengths (see Sec. 4.5). See Appendix C.2.3 for more implementation details and evaluation settings.
@@
The global watermarking results for all methods are summarized in Table 1. First, both MaskMark-D and MaskMark-ED achieve high visual fidelity, with PSNR scores above 39.5 and SSIM scores exceeding 0.98. These results outperform WAM and are only marginally lower than TrustMark and Robust-Wide. More importantly, under this high-fidelity setting, both variants of MaskMark still maintain near 100% bit accuracy, even under various valuemetric and geometric distortions. This demonstrates significantly better robustness compared to both global and local watermarking baselines. Notably, geometric distortions, which often break existing global watermarking methods, are effectively handled by our robust and reliable MaskMark framework.
@@
StegaStamp [21] 100 28.87 0.9019 0.9990 0.9976 0.6646
SepMark [23] 30 35.73 0.9876 0.9957 0.9643 0.5086
TrustMark [2] 100 41.19 0.9922 0.9996 0.9955 0.7868
EditGuard-C [28] 64 37.27 0.9332 0.9991 0.5482 0.4925
EditGuard-D [28] 64 32.30 0.8199 0.9999 0.5444 0.4975
Robust-Wide [7] 64 41.58 0.9923 1.0000 0.9944 0.4951
VINE [13] 100 36.04 0.9874 0.9997 0.9986 0.5012
@@
WAM [19] 32 39.32 0.9791 1.0000 0.9986 0.8979
MaskMark-D (Ours) 32 39.55 0.9814 1.0000 1.0000 0.9998
MaskMark-ED (Ours) 32 39.52 0.9828 1.0000 1.0000 1.0000
@@
The local watermarking results for all methods are presented in Figure 3. Global watermarking methods suffer a significant drop in extraction accuracy as the watermarked region shrinks, revealing their weakness in local watermarking tasks. In contrast, both WAM and MaskMark maintain high accuracy even with small watermarked areas, and MaskMark consistently outperforms WAM, especially under distortions, demonstrating greater robustness. Between our variants, MaskMark-ED performs better when the watermarked area is small, with the performance gap narrowing as the watermarked area increases.
@@
The visualized watermark patterns embedded by MaskMark-D and MaskMark-ED are shown in Figure 15 in Appendix D.6.
@@ 4.3 Watermark Localization Comparison
We further compare MaskMark with EditGuard [28] and WAM [19], two methods with watermark localization capabilities. Localization performance is evaluated using the local watermarking dataset described in Sec. 4.2, with Intersection-over-Union (IoU) metrics computed between the predicted and ground-truth regions for both the watermarked and unwatermarked areas.
@@ Results.
The localization results of different methods are presented in Figure 4. First, MaskMark consistently achieves the best localization performance across nearly all watermark ratios and distortion conditions, with WAM showing a noticeable performance gap and EditGuard performing significantly worse. Second, MaskMark-ED outperforms MaskMark-D in localizing small regions, whether watermarked or unwatermarked, especially under distortion conditions. Interestingly, EditGuard-C shows an unusual rise in localization performance for unwatermarked regions as the watermark area increases under non-distorted conditions, a behavior not observed in EditGuard-D, possibly due to overfitting from being trained solely on clean data. The visualized localization results of different methods are provided in Appendix D.7.
@@ 4.4 Performance Comparisons of Embedding Multiple Watermarks
We further compare MaskMark-ED with WAM [19] under a multi-watermark setting by embedding up to five distinct watermarks into separate masked regions. The evaluation considers extraction accuracy and localization performance under a stricter masking constraint (5% area per region). We report the average watermark extraction accuracy across all embedded watermarks and the mean IoU of the predicted watermark regions. Detailed setup is provided in Appendix C.2.4.
@@ Results.
Figure 5 presents the comparison results. First, our MaskMark-ED consistently outperforms WAM in both watermark extraction and localization, across all tested numbers of embedded watermarks and under various distortion conditions. Second, despite being trained solely with single-watermark supervision, MaskMark-ED generalizes well to multi-watermark settings, demonstrating strong scalability. However, under geometric distortions, MaskMark-ED’s extraction accuracy degrades as the number of embedded watermarks increases. This is mainly due to spatial transformations (particularly rotation) shrinking the watermark regions in the image corners, reducing the effective area available for extraction (see Figure 20 in Appendix D.8).
@@ 4.5 More Analysis
While WAM is constrained to 32-bit watermarks, our MaskMark readily scales to longer bit lengths. Detailed results are provided in Figure 12 in Appendix D.3. MaskMark maintains high extraction accuracy across various bit lengths. Even at 64 bits, the accuracy experiences only a slight drop and still outperforms WAM at 32 bits. While the accuracy decline becomes more noticeable at 128 bits, MaskMark consistently surpasses WAM under both no distortion and valuemetric distortion conditions. The only exception arises under geometric distortion, where a performance gap emerges when the watermarked region covers between 5% and 75%. These results underscore the scalability and optimization-friendly design of our method, emphasizing its practical advantages.
@@
While MaskMark is trained on a broad range of common distortions, it is impractical to cover all possible scenarios. Fortunately, our framework supports task-specific fine-tuning, allowing it to adapt to new and emerging threats. As a demonstration, we consider adaptive attacks based on variational autoencoders (VAEs) [10], which reconstruct images and may inadvertently remove watermark signals. To counter this, we expand the distortion pool during fine-tuning by incorporating VAE modules from Stable Diffusion v1-4 [16], Bmshj18 [1], and Cheng20 [3]. Training details are provided in Appendix C.1.1. Fine-tuning for 20k steps on a single A6000 GPU ($\sim$5 hours) significantly improves robustness against these attacks, as shown in Figure 13 in Appendix D.4, outperforming existing baselines. While there is a slight drop in robustness to other distortions, MaskMark still maintains strong overall performance. These results demonstrate the adaptability of our approach to specific threats.
@@
We compare training and inference costs of MaskMark with EditGuard [28] and WAM [19], both supporting watermark extraction and localization, summarized in Table 2. MaskMark has fewer parameters than WAM, with similar encoder and decoder sizes; WAM allocates most parameters to the decoder due to its more complex task. In training, MaskMark requires only about 1/15 of WAM’s TFLOPs. EditGuard doesn’t report training cost, but its data consumption (steps×$\times$×batch) already reaches 60% of MaskMark ’s, despite using fewer strong distortions for robustness. Fewer parameters don’t always mean lower inference overhead, as different image processing methods affect memory and speed. Although EditGuard has the fewest parameters, it uses more memory and runs slower due to embedding both localization and copyright watermarks. In contrast, MaskMark and WAM have lower latency with similar memory use.
@@
EditGuard [28] 3.6 2.6 6.2 250K 4 - - - - 0.074 s 0.080 s 0.154 s 2.15 GB
WAM [19] 1.1 96.0 97.1 3,680K 16 8 V100 90 h $1.13\times 10^{16}$ 0.015 s 0.017 s 0.032 s 2.33 GB
MaskMark (Ours) 31.1 32.2 63.3 100K 16 1 A6000 20 h $7.74\times 10^{14}$ 25.84 GB 0.009 s 0.022 s 0.031 s 2.21 GB

### v3

@@ 3.3 Inference
The inference process consists of three main stages: watermark embedding, localization, and extraction. Embedding is performed by the encoder. In MaskWM-D, the encoder embeds the watermark bits across the entire image. In MaskWM-ED, users provide both the watermark bits and a mask, allowing the encoder to embed the watermark into specific regions. Localization and extraction are performed by the decoder. As shown in Figure 2, unlike conventional global extraction methods, our approach first uses the decoder’s localization module to identify watermark-containing regions. The rest of the image is masked out by setting non-watermarked areas to zero, reducing noise during extraction. The watermark is then recovered from the retained regions. As discussed in Sec. 4.5, this strategy improves extraction robustness, especially when only small regions remain watermarked.
@@ 3.4 Usage Scenarios
This section outlines the usage scenarios of MaskWM-D and MaskWM-ED, which are tailored to different protection requirements. MaskWM-D is designed for full-image protection. It enables reliable watermark extraction even when parts of the image are tampered with, making it suitable for scenarios that demand global content integrity and tamper detection, such as copyright enforcement or image provenance verification. MaskWM-ED, on the other hand, focuses on protecting specific regions of interest, such as faces, logos, or sensitive content. It allows targeted verification or tracing if these regions are reused or misappropriated, without introducing watermarks across the entire image. This strategy not only supports privacy-aware or content-specific protection, but also tends to provide better robustness within the marked regions.
@@ 4.1 Implementation Details
For all experiments, we train MaskWM on 83k images from the MS-COCO 2014 training set (ms-coco, ) and the training details are provided in Appendix C.1.1.
@@
To ensure fair comparison, we fix the image resolution to $512\times 512$. For baseline methods that do not support this resolution, we follow the resolution scaling strategy from TrustMark (bui2023trustmark, ) to interpolate the watermark strength (see Appendix C.2.1), which has been shown to preserve watermarking performance. To ensure comparable visual fidelity across variants, we set the JND scaling factor $\mu$ to 1.3 for MaskWM-D and 1.75 for MaskWM-ED. For robustness evaluation, we separately assess valuemetric and geometric distortions. For valuemetric robustness, we randomly sample from a set of ten common distortions, including JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. For geometric robustness, we randomly sample from three typical transformations: Rotation, Perspective, and Horizontal Flip. These distortions collectively cover the vast majority of real-world transformations that watermarked images are likely to encounter in practical scenarios. Detailed parameter settings for each distortion are provided in Appendix C.2.2. We also evaluate robustness under each specific distortion individually and the detailed results are presented in Appendix D.2.
@@ 4.2 Global and Local Watermarking Comparison
We compare MaskWM with seven recent open-source watermarking methods, including both global (e.g., StegaStamp (tancik2020stegastamp, ), SepMark (wu2023sepmark, ), TrustMark (bui2023trustmark, ), EditGuard (zhang2024editguard, ), Robust-Wide (hu2024robust, ), VINE (lu2024robust, )) and local (WAM (sander2025watermark, )) approaches. We use the clean (EditGuard-C) and degraded (EditGuard-D) variants of EditGuard, and the robust version (VINE-R) of VINE. For global watermarking, we evaluate on 1k images from the MS-COCO 2014 validation set using PSNR, SSIM, and Bit Accuracy, where PSNR and SSIM measure the visual quality of watermarked images, and Bit Accuracy evaluates watermark extraction performance. For local watermarking, we use all 41k validation images and evaluate Bit Accuracy under different watermarked area ratios. By default, MaskWM embeds 32 bits for fair comparison with WAM, though MaskWM supports flexible bit lengths (see Sec. 4.5). See Appendix C.2.3 for more implementation details and evaluation settings.
@@
The global watermarking results for all methods are summarized in Table 1. First, both MaskWM-D and MaskWM-ED achieve high visual fidelity, with PSNR scores above 39.5 and SSIM scores exceeding 0.98. These results outperform WAM and are only marginally lower than TrustMark and Robust-Wide. More importantly, under this high-fidelity setting, both variants of MaskWM still maintain near 100% bit accuracy, even under various valuemetric and geometric distortions. This demonstrates significantly better robustness compared to both global and local watermarking baselines. Notably, geometric distortions, which often break existing global watermarking methods, are effectively handled by our robust and reliable MaskWM framework.
@@
StegaStamp (tancik2020stegastamp, ) 100 28.87 0.9019 0.9990 0.9976 0.6646
SepMark (wu2023sepmark, ) 30 35.73 0.9876 0.9957 0.9643 0.5086
TrustMark (bui2023trustmark, ) 100 41.19 0.9922 0.9996 0.9955 0.7868
EditGuard-C (zhang2024editguard, ) 64 37.27 0.9332 0.9991 0.5482 0.4925
EditGuard-D (zhang2024editguard, ) 64 32.30 0.8199 0.9999 0.5444 0.4975
Robust-Wide (hu2024robust, ) 64 41.58 0.9923 1.0000 0.9944 0.4951
VINE (lu2024robust, ) 100 36.04 0.9874 0.9997 0.9986 0.5012
@@
WAM (sander2025watermark, ) 32 39.32 0.9791 1.0000 0.9986 0.8979
MaskWM-D (Ours) 32 39.55 0.9814 1.0000 1.0000 0.9998
MaskWM-ED (Ours) 32 39.52 0.9828 1.0000 1.0000 1.0000
@@
The local watermarking results for all methods are presented in Figure 3. Global watermarking methods suffer a significant drop in extraction accuracy as the watermarked region shrinks, revealing their weakness in local watermarking tasks. In contrast, both WAM and MaskWM maintain high accuracy even with small watermarked areas, and MaskWM consistently outperforms WAM, especially under distortions, demonstrating greater robustness. Between our variants, MaskWM-ED performs better when the watermarked area is small, with the performance gap narrowing as the watermarked area increases.
@@
The visualized watermark patterns embedded by MaskWM-D and MaskWM-ED are shown in Figure 15 in Appendix D.6.
@@ 4.3 Watermark Localization Comparison
We further compare MaskWM with EditGuard (zhang2024editguard, ) and WAM (sander2025watermark, ), two methods with watermark localization capabilities. Localization performance is evaluated using the local watermarking dataset described in Sec. 4.2, with Intersection-over-Union (IoU) metrics computed between the predicted and ground-truth regions for both the watermarked and unwatermarked areas.
@@ Results.
The localization results of different methods are presented in Figure 4. First, MaskWM consistently achieves the best localization performance across nearly all watermark ratios and distortion conditions, with WAM showing a noticeable performance gap and EditGuard performing significantly worse. Second, MaskWM-ED outperforms MaskWM-D in localizing small regions, whether watermarked or unwatermarked, especially under distortion conditions. Interestingly, EditGuard-C shows an unusual rise in localization performance for unwatermarked regions as the watermark area increases under non-distorted conditions, a behavior not observed in EditGuard-D, possibly due to overfitting from being trained solely on clean data. The visualized localization results of different methods are provided in Appendix D.7.
@@ 4.4 Performance Comparisons of Embedding Multiple Watermarks
We further compare MaskWM-ED with WAM (sander2025watermark, ) under a multi-watermark setting by embedding up to five distinct watermarks into separate masked regions. The evaluation considers extraction accuracy and localization performance under a stricter masking constraint (5% area per region). We report the average watermark extraction accuracy across all embedded watermarks and the mean IoU of the predicted watermark regions. Detailed setup is provided in Appendix C.2.4.
@@ Results.
Figure 5 presents the comparison results. First, our MaskWM-ED consistently outperforms WAM in both watermark extraction and localization, across all tested numbers of embedded watermarks and under various distortion conditions. Second, despite being trained solely with single-watermark supervision, MaskWM-ED generalizes well to multi-watermark settings, demonstrating strong scalability. However, under geometric distortions, MaskWM-ED’s extraction accuracy degrades as the number of embedded watermarks increases. This is mainly due to spatial transformations (particularly rotation) shrinking the watermark regions in the image corners, reducing the effective area available for extraction (see Figure 20 in Appendix D.8).
@@ 4.5 More Analysis
While WAM is constrained to 32-bit watermarks, our MaskWM readily scales to longer bit lengths. Detailed results are provided in Figure 12 in Appendix D.3. MaskWM maintains high extraction accuracy across various bit lengths. Even at 64 bits, the accuracy experiences only a slight drop and still outperforms WAM at 32 bits. While the accuracy decline becomes more noticeable at 128 bits, MaskWM consistently surpasses WAM under both no distortion and valuemetric distortion conditions. The only exception arises under geometric distortion, where a performance gap emerges when the watermarked region covers between 5% and 75%. These results underscore the scalability and optimization-friendly design of our method, emphasizing its practical advantages.
@@
While MaskWM is trained on a broad range of common distortions, it is impractical to cover all possible scenarios. Fortunately, our framework supports task-specific fine-tuning, allowing it to adapt to new and emerging threats. As a demonstration, we consider adaptive attacks based on variational autoencoders (VAEs) (kingma2014auto, ), which reconstruct images and may inadvertently remove watermark signals. To counter this, we expand the distortion pool during fine-tuning by incorporating VAE modules from Stable Diffusion v1-4 (rombach2022high, ), Bmshj18 (ballé2018variational, ), and Cheng20 (cheng2020learned, ). Training details are provided in Appendix C.1.1. Fine-tuning for 20k steps on a single A6000 GPU ($\sim$5 hours) significantly improves robustness against these attacks, as shown in Figure 13 in Appendix D.4, outperforming existing baselines. While there is a slight drop in robustness to other distortions, MaskWM still maintains strong overall performance. These results demonstrate the adaptability of our approach to specific threats.
@@
[…]
@@
EditGuard (zhang2024editguard, ) 3.6 2.6 6.2 250K 4 - - - - 0.074 s 0.080 s 0.154 s 2.15 GB
WAM (sander2025watermark, ) 1.1 96.0 97.1 3,680K 16 8 V100 90 h $1.13\times 10^{16}$ - 0.015 s 0.017 s 0.032 s 2.33 GB
MaskWM (Ours) 31.1 32.2 63.3 100K 16 1 A6000 20 h $7.74\times 10^{14}$ 25.84 GB 0.009 s 0.022 s 0.031 s 2.21 GB